OpenStudy (anonymous):

write an equation in slope intercept form for the line that contains the given points, (-8,6) and (-6,9) write an equation in slope intercept form for the line that contains the given points, (-8,6) and (-6,9) @Mathematics

OpenStudy (anonymous):

First you have to find the slope. m = (y1-y2)/(x1-x2) m=(9-6)/(-6 - (-8)) m= 3/2 Now you take the equation y = 3/2 x +b, pick one of the points, plug in the x and y values, and solve for b.

OpenStudy (anonymous):

For example if we use (-8,6) we get 6 = (3/2) (-8) + b 6 = - 12 +b (add 12 to both sides) b = 18 so y = 3/2 x + 18